OpenStudy (anonymous):

convert 0.044% to a decimal....it appears to be in a decimal..unless i put it as 44.0 % plz help

OpenStudy (anonymous):

divide it by 100

OpenStudy (slaaibak):

Nope.To put a percentage in a decimal, you divide by 100

OpenStudy (anonymous):

thx =)

OpenStudy (turingtest):

so what answer did you get?

OpenStudy (anonymous):

hang on

OpenStudy (anonymous):

4.4% is that right?

OpenStudy (anonymous):

it threw me off because of it already being in decimal form

OpenStudy (turingtest):

not quite... whatever the percentage is, divide by 100 to get it as a decimal. It doesn't matter if it's a decimal already. i.e. 54%=54/100=0.54 6%=6/100=0.06 0.25%=0.25/100=0.0025 0.007%=0.007/100=0.00007 so what is 0.044%=?

OpenStudy (anonymous):

2272.72 is all i came up with i divided each way by 100 an these were my answers

OpenStudy (turingtest):

you divided 0.044/100 and got that? o-0 try again \[0.044\%=\frac{0.044}{100}=?\]

OpenStudy (anonymous):

that or 4.4 -04 r the 2 answers i got

OpenStudy (anonymous):

idk how it seems stupid

OpenStudy (turingtest):

Here's the answer. I only give one away for free, just so you know, so try to figure out what you did wrong because next time I won't give it away so easily\[0.044\%=\frac{0.044}{100}=0.00044\]just divide the percentage by 100... a job for your calculator if that gives you trouble...

OpenStudy (anonymous):

ok thx but i have to know how to figure it out for myself i dont think you give me the answer so eaSILY but i do thank u.....now on the real side i am telling you my calculator gave me the 2 answers i presented to you so im confused unless to the decimal you add extra zeros???

OpenStudy (anonymous):

to divide a decimal by 100 move the decimal point two places to the LEFT - if there are no '0's insert them

OpenStudy (anonymous):

thx jimmy thats what i was trying to say above

OpenStudy (turingtest):

I put 0.044/100 into my calculator and it says 4.4E-4 is that what you see?

OpenStudy (anonymous):